OpenStudy (anonymous):

if csc(theta)=10 what are the values of cos(theta) and cot(theta)

OpenStudy (anonymous):

\[cosec \theta =10 \rightarrow \sin \theta =\frac{1}{cosec \theta}\] \[\sin \theta =\frac{1}{10}\] \[\cos^2 \theta = 1-\sin^2 \theta \] \[= 1- (\frac{1}{10})^2\] \[=1- \frac{1}{100} = \frac{100-1}{100}=\frac{99}{100}\] \[\cos \theta = \sqrt{0.99}=0.9949\]